OpenStudy (anonymous):

Evaluate g(x) = -8x^3 + 5x^2 + 14; x = -3 -275 275 -230 230

OpenStudy (mertsj):

\[-8(-3)^3+5(-3)^2+14=-8(-27)+5(9)+14=216+45+14\]

OpenStudy (anonymous):

275 is the answer just substitute -3 wherever you see an x and do the resulting sum